What is the speed of a plane traveling 1.530 kilometers in 2 hours and 45 minutes?

Answer by stanbon(75887)

What is the speed of a plane traveling 1.530 kilometers in 2 hours and 45 minutes?
--------------------
Note: 2 hr 45 min = 2 3/4 = (11/4)hr
--------------------
rate = distance/time
---
rate = 1.53 km/(11/4)hr = 0.5564 km/hr
=========================================
Cheers,
Stan H.
===========